MalayaJanae:

Evaluate the following expression when x = 4 and y = 5:

supie:

What is the expression?

MalayaJanae:

Fraction with x squared plus y cubed in the numerator and 2 plus x in the denominator.

supie:

Can you take a screenshot please?

MalayaJanae:

Chromebook. im new here..

supie:

Oh ok Well does the expression look like this, \(\LARGE\frac{x^2+y^3}{2+x}\)?

supie:

@MalayaJanae

supie:

So we basically just plug in the numbers into \(\LARGE\frac{x^2+y^3}{2+x}\) x=4 y=5 So we plug in `4` for every x and `5` for every y So we get \(\LARGE\frac{4^2+5^3}{2+4}\)